OpenStudy (anonymous):

Find the area of a circle with an area of pi(16^2 + 24x + 9)

OpenStudy (jdoe0001):

"Find the \(\bf area\) of a circle with an \(\bf area\) of pi(16^2 + 24x + 9)" ^ found, is right here

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

16x+9?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well, the area is already a given it seems

OpenStudy (anonymous):

Sorry haha, I didn't realize what I wrote

OpenStudy (anonymous):

Let me rephrase it

OpenStudy (anonymous):

Find the radius of a circle with an area of \[\pi(16^2+24x+9)\]

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ \textit{area of a circle}=\pi r^2}} \\ \quad \\ \quad \\ Area=\pi(16^2+24x+9)\implies {\color{blue}{ \pi r^2}}=\pi(16^2+24x+9)\) solve for "r", the value would end up in terms of "x"